Let X~Bin(n,p).Find E(e^tx) where t is a constant
1
Expert's answer
2020-10-05T17:39:28-0400

Using independence, EetΣXi=E"\Pi" etXi="\Pi"EetXi=(pet+(1-p))n,

where the Xi are independent Bernoulli random variables. Equivalently

"\sum^n_{k=0}e^{tk}(^n_k)p^k(1-p)^{n-k}=\sum^n_{k=0}(^n_k)(pe^{t})^k(1-p)^{n-k}=(pe^t+(1-p)^n)" by the Binomial formula.